OpenStudy (anonymous):

A plane travels from Orlando to Denver & back again. On the 5-hr trip from Orlando to Denver, the plane has a tailwind of 40 mph. On the return trip from Denver to Orlando, the plane faces a headwind of 40 mph. This trip takes 6 hrs. What is the speed of the airplane in still air?

OpenStudy (anonymous):

5(x+40) = 6(x-40) 5x+200=6x-240 6x-5x=240+200 x=440 mph Is this correct?